A horizontal rifle shoots a bullet 100 m to a target. During flight, the bullet drops 6.5 cm. How fast was the bullet traveling?

a) 767 m/s

b) 802 m/s

c) 275 m/s

d) 868 m/s

Which of the following is an example of the type of force that acts at a distance?

a) all of the answers

b) gravitational

c) magnetic

d) electric
Answer #1

here

speed = distance / time

t = 100 / v ...........(1)

then by using the second equation of motion

y = ut + 0.5 * a * t^2

0.065 = 0 + 0.5 * 9.8 * t^2

0.065 = 4.9 t^2 ...............(2)

then put the value of t form equation (1) to equation(2)

0.065 = 4.9 * (100 / v)^2

v = 868 m /s

so the correct answer is d) 868 m/s

a) all of the answers